# Bo Du1† Xuekang Zhu1,2† Xiaochen Ma3† Chenfan Qu4, 2† Kaiwen Feng1† Zhe Yang1 Chi-Man Pun5 Jian Liu2‡ Jizhe Zhou1‡

### v2

@@ Results.
Table 3 in green background presents the AUC scores for image-level classification for AIGC benchmark, divided into in-domain results on the test set split of DiffusionForensics wang2023dire , and cross-domain results on different generative models and the total set on GenImage zhu2023genimage . The results show that AIGC SoTAs generally achieve excellent performance on the DiffusionForensics test set, which shares the same source as the training set, and also perform well on datasets composed of diffusion-based generated images like ADM, VQDM, and GLIDE that are similar to the training data. However, the relatively poor generalization to generative models like Midjourney and Wukong highlights areas for improvement and provides guidance for future model development.
@@ 4.2 Document Image Manipulation Localization Benchmark
DualNet dualnet (APSIPA23) 1.0000 0.9986 0.9172 0.8495 0.9901 0.9835 0.8127 0.8081 0.7925 0.8902
HiFiNet HiFi-Net2023 (CVPR23) 1.0000 0.9998 0.8407 0.7210 0.9999 0.9991 0.6765 0.6747 0.6675 0.8160
Synthbuster bammey2023synthbuster (OJSP23) 0.6662 0.6226 0.8550 0.4565 0.6308 0.7719 0.4606 0.4614 0.3902 0.5762
UnivFD univfd (CVPR23) 0.9947 0.8400 0.9687 0.5427 0.9663 0.9541 0.7078 0.7072 0.6781 0.7920
MVSS-Net MVSS_2021 (ICCV21) 0.9992 0.9706 0.9520 0.6242 0.8894 0.9380 0.7207 0.7192 0.6089 0.7997
CAT-Net CAT-Net2022 (IJCV22) 0.9985 0.9391 0.9327 0.6423 0.8226 0.7987 0.7171 0.7142 0.6156 0.7707
Trufor trufor2023 (CVPR23) 0.9999 0.9750 0.9316 0.7079 0.9340 0.9144 0.8683 0.8687 0.8018 0.8752
IML-ViT ma2023iml (Arxiv) 1.0000 0.9594 0.9152 0.8086 0.9577 0.9464 0.8924 0.8877 0.8006 0.8957
Mesorch zhu2025mesoscopic (AAAI25) 0.9998 0.9754 0.9667 0.7011 0.9331 0.9253 0.8168 0.8179 0.7441 0.8582
@@
Existing datasets for document image manipulation localization can be broadly categorized into two types: high-fidelity non-sliced datasets, including T-SROIE wang2022tsroie , OSTF qu2025ostf , TPIC-13 wang2022tpic , and RTM luo2025rtm ; and sliced datasets, represented by Doctamper qu2023doctamper . The primary difference lies in whether the images are preprocessed using patch-wise slicing.
@@
We employ two open-source models, DTD qu2023doctamper and FFDN chen2024ffdn , and reproduce two closed-source models, CATFB song2025caftb and TIFDM dong2024tifdm . All details are provided in Appendix D.2.4 and E.2.
@@ Results.
Following the original protocols qu2023doctamper ; chen2024ffdn , each detector is trained on its designated training split and evaluated on the corresponding test split. As shown in Table 4, three models consistently achieve top performance: FFDN and DTD, both designed specifically for document forensics, and Cat-Net, an IMDL-based model. Notably, all three methods incorporate JPEG-specific priors, such as DCT coefficients and quantization tables, highlighting the discriminative value of compression artifacts for manipulation localization in document images.
@@
CAFTB song2025caftb (TOMM24) 0.2917 0.3770 0.3275 0.3321 0.9165 0.6478 0.8394 0.2493 0.5213
DTD qu2023doctamper (CVPR23) 0.6856 0.7392 0.8031 0.7426 0.9205 0.5626 0.8341 0.1720 0.6739
FFDN chen2024ffdn (ECCV24) 0.8773 0.7392 0.8212 0.8126 0.9137 0.5586 0.8186 0.1589 0.6982
TIFDM dong2024tifdm (TCE24) 0.0896 0.2572 0.2585 0.2018 0.8942 0.5410 0.7972 0.0591 0.4138
MVSS-Net MVSS_2021 (ICCV21) 0.2066 0.3710 0.3810 0.3195 0.8756 0.5254 0.7864 0.1314 0.4682
PSCC-Net liu2022pscc (TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.9305 0.5697 0.7894 0.1971 0.5375
Cat-Net CAT-Net2022 (IJCV22) 0.7600 0.6405 0.7644 0.7216 0.8726 0.5371 0.7947 0.1174 0.6410
IML-ViT ma2023iml (Arxiv) 0.4688 0.5117 0.4486 0.4764 0.8731 0.5128 0.7202 0.0920 0.5182
Trufor trufor2023 (CVPR23) 0.2613 0.3124 0.2517 0.2751 0.9095 0.6485 0.8197 0.2302 0.4905
Mesorch zhu2025mesoscopic (AAAI25) 0.4231 0.4586 0.4651 0.4489 0.9049 0.4804 0.7607 0.1314 0.5177
@@
CAFTB song2025caftb (TOMM24) 0.2917 0.3770 0.3275 0.3321 0.2617 0.1194 0.3007 0.0328 0.1787 0.2444
DTD qu2023doctamper (CVPR23) 0.6856 0.7392 0.8031 0.7426 0.5245 0.1241 0.2835 0.0575 0.2474 0.4596
FFDN chen2024ffdn (ECCV24) 0.8773 0.7392 0.8212 0.8126 0.5330 0.2409 0.3572 0.0708 0.3005 0.5199
TIFDM dong2024tifdm (TCE24) 0.0896 0.2572 0.2585 0.2018 0.0582 0.0058 0.0134 0.0176 0.0238 0.1000
MVSS-Net MVSS_2021 (ICCV23) 0.2066 0.3710 0.3810 0.3195 0.1870 0.0373 0.1134 0.0268 0.0911 0.1890
PSCC-Net liu2022pscc (TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.5168 0.4414 0.5495 0.1255 0.4083 0.4156
Cat-Net CAT-Net2022 (IJCV22) 0.7600 0.6405 0.7644 0.7216 0.6085 0.1777 0.3430 0.0630 0.2981 0.4796
IML-ViT ma2023iml (Arxiv) 0.4688 0.5117 0.4486 0.4764 0.4269 0.2101 0.2563 0.0764 0.2424 0.3427
Trufor trufor2023 (CVPR23) 0.2613 0.3124 0.2517 0.2751 0.2126 0.0464 0.1038 0.0342 0.0993 0.1746
Mesorch zhu2025mesoscopic (AAAI25) 0.4231 0.4586 0.4651 0.4489 0.2937 0.1388 0.2408 0.0405 0.1785 0.2944
@@ 5 Image Forensic Fusion Protocol
To explore the performance of different models under a unified forensic protocol, we implement an image forensic fusion protocol (IFF-Protocol) inspired by CAT-Net’s training data construction strategy. The IFF-Protocol defines the training set as a combination of Deepfake, IMDL, AIGC, and Document data, where each training epoch samples an equal number of instances from each domain at random. During training, we select FaceForensics++ rossler2019faceforensics++ from Deepfake, CASIAv2 CASIA_2013 from IMDL, GenImage zhu2023genimage from AIGC, and OSTF qu2025ostf , RealTextManipulation luo2025rtm , T-SROIE wang2022tsroie , and Tampered-IC13 wang2022tpic from the Document. We use the smallest dataset, CASIAv2 with 12,641 samples, as the sampling number for each epoch. During testing, we evaluate the models directly on datasets from different domains without fine-tuning.
@@
Resnet Resnet_2016 $44.55$ $20.59$ Capsule-Net nguyen2019capsule $3.90$ $32.43$ Mesorch zhu2025mesoscopic $85.75$ $57.95$
Xception chollet2017xception $22.86$ $12.05$ RECCE cao2022recce $25.83$ $16.18$ DualNet dualnet $7.99$ $66.34$
EfficientNet tan2019efficientnet $19.34$ $4.14$ SPSL liu2021spsl $20.81$ $12.06$ HiFiNet HiFi-Net2023 $6.89$ $145.00$
ConvNeXt Convnet_2022 $50.22$ $22.74$ MVSS-Net MVSS_2021 $147.00$ $83.08$ UnivFD univfd $428.00$ $156.00$
Segformer SegFormer_2021 $44.07$ $16.29$ Trufor trufor2023 $68.70$ $112.00$ DTD qu2023doctamper $67.07$ $272.00$
Swin Swin_2021 $49.61$ $28.65$ IML-ViT ma2023iml $91.78$ $80.37$ FFDN chen2024ffdn $89.20$ $453.00$
@@
Resnet Resnet_2016 (CVPR16) 0.681 0.730 0.793 0.482 0.533 0.738 0.619 0.797 0.951 0.681 0.662 0.697
Xception chollet2017xception (CVPR17) 0.728 0.719 0.870 0.465 0.537 0.756 0.757 0.980 0.966 0.762 0.734 0.752
EfficientNet tan2019efficientnet (ICML19) 0.504 0.535 0.517 0.623 0.506 0.483 0.544 0.597 0.884 0.581 0.512 0.571
Segformer SegFormer_2021 (NIPS21) 0.691 0.748 0.862 0.409 0.562 0.824 0.805 0.998 0.980 0.866 0.736 0.771
Swin Swin_2021 (ICCV21) 0.771 0.746 0.901 0.636 0.631 0.864 0.915 0.999 0.990 0.856 0.758 0.824
ConvNext Convnet_2022 (CVPR22) 0.794 0.784 0.911 0.625 0.598 0.825 0.895 1.000 0.994 0.849 0.762 0.822
Capsule-Net nguyen2019capsule (ICASSP19) 0.613 0.660 0.699 0.330 0.527 0.745 0.546 0.971 0.946 0.704 0.670 0.674
RECCE cao2022recce (CVPR22) 0.634 0.602 0.727 0.506 0.492 0.642 0.684 0.906 0.542 0.688 0.555 0.634
SPSL liu2021spsl (CVPR21) 0.730 0.726 0.876 0.419 0.545 0.759 0.770 0.987 0.972 0.769 0.738 0.754
Sia sia (ECCV22) 0.629 0.584 0.671 0.653 0.483 0.626 0.593 0.748 0.610 0.677 0.574 0.622
Effort effort (ICML25) 0.805 0.846 0.930 0.979 0.861 0.943 0.930 0.992 0.960 0.834 0.732 0.892
MVSS-Net MVSS_2021 (ICCV21) 0.713 0.700 0.857 0.298 0.539 0.795 0.671 0.994 0.978 0.806 0.741 0.736
Trufor trufor2023 (CVPR23) 0.642 0.698 0.832 0.306 0.564 0.808 0.726 0.996 0.979 0.805 0.732 0.735
IML-ViT ma2023iml (Arxiv) 0.750 0.726 0.851 0.483 0.556 0.819 0.627 0.991 0.972 0.800 0.703 0.753
Mesorch zhu2025mesoscopic (AAAI25) 0.767 0.814 0.867 0.285 0.570 0.773 0.629 0.996 0.982 0.819 0.739 0.749
DualNet dualnet (APSIPA23) 0.637 0.552 0.540 0.268 0.517 0.748 0.899 0.988 0.935 0.658 0.657 0.673
HiFiNet HiFi-Net2023 (CVPR23) 0.587 0.611 0.648 0.745 0.534 0.677 0.575 0.756 0.937 0.663 0.615 0.668
UnivFD univfd (CVPR23) 0.690 0.671 0.798 0.886 0.786 0.785 0.742 0.813 0.938 0.684 0.569 0.760
FatFormer fatformer (CVPR24) 0.842 0.770 0.866 0.199 0.585 0.784 0.941 0.999 0.983 0.806 0.751 0.758
CO-SPY co-spy (CVPR25) 0.819 0.780 0.875 0.460 0.716 0.779 0.940 0.989 0.969 0.836 0.748 0.829
DTD qu2023doctamper (CVPR23) 0.498 0.520 0.490 0.679 0.498 0.506 0.457 0.499 0.748 0.595 0.496 0.544
FFDN chen2024ffdn (ECCV24) 0.714 0.699 0.871 0.553 0.624 0.927 0.999 1.000 0.997 0.893 0.782 0.824
@@
Table 7 shows the AUC scores of backbones and domain-specific SoTA methods on datasets from four domains under the IFF-Protocol, in which DFD refers to DeepFakeDetection dfd2019 , DF refers to DiffusionForensics wang2023dire , and RTM refers to RealTextManipulation luo2025rtm . We provide detailed results in Appendix F.
The results show that surprisingly, visual backbones such as ConvNeXt Convnet_2022 and Swin Transformer Swin_2021 outperform almost all domain-specific SoTA methods, indicating that backbones demonstrate greater potential when trained on more unified fake images. Meanwhile, domain-specific SoTAs do not necessarily retain their superiority within their own tasks. For instance, UnivFD univfd , a CLIP-based fine-tuned model for AIGC detection, demonstrates strong performance on the IMD2020 IMD20_2020 from IMDL, revealing valuable insights into the transferability of cross-task methods.

### v3

@@ Results.
Table 3 in green background presents the AUC scores for image-level classification for AIGC benchmark, divided into in-domain results on the test set split of DiffusionForensics Wang et al. (2023), and cross-domain results on different generative models and the total set on GenImage Zhu et al. (2023). The results show that AIGC SoTAs generally achieve excellent performance on the DiffusionForensics test set, which shares the same source as the training set, and also perform well on datasets composed of diffusion-based generated images like ADM, VQDM, and GLIDE that are similar to the training data. However, the relatively poor generalization to generative models like Midjourney and Wukong highlights areas for improvement and provides guidance for future model development.
@@ 4.2 Document Image Manipulation Localization Benchmark
DualNet Xi et al. (2023)(APSIPA23) 1.0000 0.9986 0.9172 0.8495 0.9901 0.9835 0.8127 0.8081 0.7925 0.8902
HiFiNet Guo et al. (2023)(CVPR23) 1.0000 0.9998 0.8407 0.7210 0.9999 0.9991 0.6765 0.6747 0.6675 0.8160
Synthbuster Bammey (2023)(OJSP23) 0.6662 0.6226 0.8550 0.4565 0.6308 0.7719 0.4606 0.4614 0.3902 0.5762
UnivFD Ojha et al. (2023)(CVPR23) 0.9947 0.8400 0.9687 0.5427 0.9663 0.9541 0.7078 0.7072 0.6781 0.7920
MVSS-Net Chen et al. (2021)(ICCV21) 0.9992 0.9706 0.9520 0.6242 0.8894 0.9380 0.7207 0.7192 0.6089 0.7997
CAT-Net Kwon et al. (2022)(IJCV22) 0.9985 0.9391 0.9327 0.6423 0.8226 0.7987 0.7171 0.7142 0.6156 0.7707
Trufor Guillaro et al. (2023)(CVPR23) 0.9999 0.9750 0.9316 0.7079 0.9340 0.9144 0.8683 0.8687 0.8018 0.8752
IML-ViT Ma et al. (2023b)(Arxiv) 1.0000 0.9594 0.9152 0.8086 0.9577 0.9464 0.8924 0.8877 0.8006 0.8957
Mesorch Zhu et al. (2025)(AAAI25) 0.9998 0.9754 0.9667 0.7011 0.9331 0.9253 0.8168 0.8179 0.7441 0.8582
@@
Existing datasets for document image manipulation localization can be broadly categorized into two types: high-fidelity non-sliced datasets, including T-SROIE Wang et al. (2022b), OSTF Qu et al. (2025), TPIC-13 Wang et al. (2022a), and RTM Luo et al. (2025); and sliced datasets, represented by Doctamper Qu et al. (2023). The primary difference lies in whether the images are preprocessed using patch-wise slicing.
@@
We employ two open-source models, DTD Qu et al. (2023) and FFDN Chen et al. (2024), and reproduce two closed-source models, CATFB Song et al. (2025) and TIFDM Dong et al. (2024). All details are provided in Appendix D.2.4 and E.2.
@@ Results.
Following the original protocols Qu et al. (2023); Chen et al. (2024), each detector is trained on its designated training split and evaluated on the corresponding test split. As shown in Table 4, three models consistently achieve top performance: FFDN and DTD, both designed specifically for document forensics, and Cat-Net, an IMDL-based model. Notably, all three methods incorporate JPEG-specific priors, such as DCT coefficients and quantization tables, highlighting the discriminative value of compression artifacts for manipulation localization in document images.
@@
CAFTB Song et al. (2025)(TOMM24) 0.2917 0.3770 0.3275 0.3321 0.9165 0.6478 0.8394 0.2493 0.5213
DTD Qu et al. (2023)(CVPR23) 0.6856 0.7392 0.8031 0.7426 0.9205 0.5626 0.8341 0.1720 0.6739
FFDN Chen et al. (2024)(ECCV24) 0.8773 0.7392 0.8212 0.8126 0.9137 0.5586 0.8186 0.1589 0.6982
TIFDM Dong et al. (2024)(TCE24) 0.0896 0.2572 0.2585 0.2018 0.8942 0.5410 0.7972 0.0591 0.4138
MVSS-Net Chen et al. (2021)(ICCV21) 0.2066 0.3710 0.3810 0.3195 0.8756 0.5254 0.7864 0.1314 0.4682
PSCC-Net Liu et al. (2022a)(TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.9305 0.5697 0.7894 0.1971 0.5375
Cat-Net Kwon et al. (2022)(IJCV22) 0.7600 0.6405 0.7644 0.7216 0.8726 0.5371 0.7947 0.1174 0.6410
IML-ViT Ma et al. (2023b)(Arxiv) 0.4688 0.5117 0.4486 0.4764 0.8731 0.5128 0.7202 0.0920 0.5182
Trufor Guillaro et al. (2023)(CVPR23) 0.2613 0.3124 0.2517 0.2751 0.9095 0.6485 0.8197 0.2302 0.4905
Mesorch Zhu et al. (2025)(AAAI25) 0.4231 0.4586 0.4651 0.4489 0.9049 0.4804 0.7607 0.1314 0.5177
@@
CAFTB Song et al. (2025)(TOMM24) 0.2917 0.3770 0.3275 0.3321 0.2617 0.1194 0.3007 0.0328 0.1787 0.2444
DTD Qu et al. (2023)(CVPR23) 0.6856 0.7392 0.8031 0.7426 0.5245 0.1241 0.2835 0.0575 0.2474 0.4596
FFDN Chen et al. (2024)(ECCV24) 0.8773 0.7392 0.8212 0.8126 0.5330 0.2409 0.3572 0.0708 0.3005 0.5199
TIFDM Dong et al. (2024)(TCE24) 0.0896 0.2572 0.2585 0.2018 0.0582 0.0058 0.0134 0.0176 0.0238 0.1000
MVSS-Net Chen et al. (2021)(ICCV23) 0.2066 0.3710 0.3810 0.3195 0.1870 0.0373 0.1134 0.0268 0.0911 0.1890
PSCC-Net Liu et al. (2022a)(TCSVT22) 0.3855 0.3931 0.4972 0.4253 0.5168 0.4414 0.5495 0.1255 0.4083 0.4156
Cat-Net Kwon et al. (2022)(IJCV22) 0.7600 0.6405 0.7644 0.7216 0.6085 0.1777 0.3430 0.0630 0.2981 0.4796
IML-ViT Ma et al. (2023b)(Arxiv) 0.4688 0.5117 0.4486 0.4764 0.4269 0.2101 0.2563 0.0764 0.2424 0.3427
Trufor Guillaro et al. (2023)(CVPR23) 0.2613 0.3124 0.2517 0.2751 0.2126 0.0464 0.1038 0.0342 0.0993 0.1746
Mesorch Zhu et al. (2025)(AAAI25) 0.4231 0.4586 0.4651 0.4489 0.2937 0.1388 0.2408 0.0405 0.1785 0.2944
@@ 5 Image Forensic Fusion Protocol
To explore the performance of different models under a unified forensic protocol, we implement an image forensic fusion protocol (IFF-Protocol) inspired by CAT-Net’s training data construction strategy. The IFF-Protocol defines the training set as a combination of Deepfake, IMDL, AIGC, and Document data, where each training epoch samples an equal number of instances from each domain at random. During training, we select FaceForensics++ Rossler et al. (2019) from Deepfake, CASIAv2 Dong et al. (2013) from IMDL, GenImage Zhu et al. (2023) from AIGC, and OSTF Qu et al. (2025), RealTextManipulation Luo et al. (2025), T-SROIE Wang et al. (2022b), and Tampered-IC13 Wang et al. (2022a) from the Document. We use the smallest dataset, CASIAv2 with 12,641 samples, as the sampling number for each epoch. During testing, we evaluate the models directly on datasets from different domains without fine-tuning.
@@
Resnet He et al. (2016) $44.55$ $20.59$ Capsule-Net Nguyen et al. (2019) $3.90$ $32.43$ Mesorch Zhu et al. (2025) $85.75$ $57.95$
Xception Chollet (2017) $22.86$ $12.05$ RECCE Cao et al. (2022) $25.83$ $16.18$ DualNet Xi et al. (2023) $7.99$ $66.34$
EfficientNet Tan and Le (2019) $19.34$ $4.14$ SPSL Liu et al. (2021a) $20.81$ $12.06$ HiFiNet Guo et al. (2023) $6.89$ $145.00$
ConvNeXt Liu et al. (2022b) $50.22$ $22.74$ MVSS-Net Chen et al. (2021) $147.00$ $83.08$ UnivFD Ojha et al. (2023) $428.00$ $156.00$
Segformer Xie et al. (2021) $44.07$ $16.29$ Trufor Guillaro et al. (2023) $68.70$ $112.00$ DTD Qu et al. (2023) $67.07$ $272.00$
Swin Liu et al. (2021b) $49.61$ $28.65$ IML-ViT Ma et al. (2023b) $91.78$ $80.37$ FFDN Chen et al. (2024) $89.20$ $453.00$
@@
Resnet He et al. (2016)(CVPR16) 0.681 0.730 0.793 0.482 0.533 0.738 0.619 0.797 0.951 0.681 0.662 0.697
Xception Chollet (2017)(CVPR17) 0.728 0.719 0.870 0.465 0.537 0.756 0.757 0.980 0.966 0.762 0.734 0.752
EfficientNet Tan and Le (2019)(ICML19) 0.504 0.535 0.517 0.623 0.506 0.483 0.544 0.597 0.884 0.581 0.512 0.571
Segformer Xie et al. (2021)(NIPS21) 0.691 0.748 0.862 0.409 0.562 0.824 0.805 0.998 0.980 0.866 0.736 0.771
Swin Liu et al. (2021b)(ICCV21) 0.771 0.746 0.901 0.636 0.631 0.864 0.915 0.999 0.990 0.856 0.758 0.824
ConvNext Liu et al. (2022b)(CVPR22) 0.794 0.784 0.911 0.625 0.598 0.825 0.895 1.000 0.994 0.849 0.762 0.822
Capsule-Net Nguyen et al. (2019)(ICASSP19) 0.613 0.660 0.699 0.330 0.527 0.745 0.546 0.971 0.946 0.704 0.670 0.674
RECCE Cao et al. (2022)(CVPR22) 0.634 0.602 0.727 0.506 0.492 0.642 0.684 0.906 0.542 0.688 0.555 0.634
SPSL Liu et al. (2021a)(CVPR21) 0.730 0.726 0.876 0.419 0.545 0.759 0.770 0.987 0.972 0.769 0.738 0.754
Sia Sun et al. (2022a)(ECCV22) 0.629 0.584 0.671 0.653 0.483 0.626 0.593 0.748 0.610 0.677 0.574 0.622
Effort Yan et al. (2024c)(ICML25) 0.805 0.846 0.930 0.979 0.861 0.943 0.930 0.992 0.960 0.834 0.732 0.892
MVSS-Net Chen et al. (2021)(ICCV21) 0.713 0.700 0.857 0.298 0.539 0.795 0.671 0.994 0.978 0.806 0.741 0.736
Trufor Guillaro et al. (2023)(CVPR23) 0.642 0.698 0.832 0.306 0.564 0.808 0.726 0.996 0.979 0.805 0.732 0.735
IML-ViT Ma et al. (2023b)(Arxiv) 0.750 0.726 0.851 0.483 0.556 0.819 0.627 0.991 0.972 0.800 0.703 0.753
Mesorch Zhu et al. (2025)(AAAI25) 0.767 0.814 0.867 0.285 0.570 0.773 0.629 0.996 0.982 0.819 0.739 0.749
DualNet Xi et al. (2023)(APSIPA23) 0.637 0.552 0.540 0.268 0.517 0.748 0.899 0.988 0.935 0.658 0.657 0.673
HiFiNet Guo et al. (2023)(CVPR23) 0.587 0.611 0.648 0.745 0.534 0.677 0.575 0.756 0.937 0.663 0.615 0.668
UnivFD Ojha et al. (2023)(CVPR23) 0.690 0.671 0.798 0.886 0.786 0.785 0.742 0.813 0.938 0.684 0.569 0.760
FatFormer Liu et al. (2024)(CVPR24) 0.842 0.770 0.866 0.199 0.585 0.784 0.941 0.999 0.983 0.806 0.751 0.758
CO-SPY Cheng et al. (2025)(CVPR25) 0.819 0.780 0.875 0.460 0.716 0.779 0.940 0.989 0.969 0.836 0.748 0.829
DTD Qu et al. (2023)(CVPR23) 0.498 0.520 0.490 0.679 0.498 0.506 0.457 0.499 0.748 0.595 0.496 0.544
FFDN Chen et al. (2024)(ECCV24) 0.714 0.699 0.871 0.553 0.624 0.927 0.999 1.000 0.997 0.893 0.782 0.824
@@
Table 7 shows the AUC scores of backbones and domain-specific SoTA methods on datasets from four domains under the IFF-Protocol, in which DFD refers to DeepFakeDetection Google AI Blog (2019), DF refers to DiffusionForensics Wang et al. (2023), and RTM refers to RealTextManipulation Luo et al. (2025). We provide detailed results in Appendix F.
The results show that surprisingly, visual backbones such as ConvNeXt Liu et al. (2022b) and Swin Transformer Liu et al. (2021b) outperform almost all domain-specific SoTA methods, indicating that backbones demonstrate greater potential when trained on more unified fake images. Meanwhile, domain-specific SoTAs do not necessarily retain their superiority within their own tasks. For instance, UnivFD Ojha et al. (2023), a CLIP-based fine-tuned model for AIGC detection, demonstrates strong performance on the IMD2020 Novozamsky et al. (2020) from IMDL, revealing valuable insights into the transferability of cross-task methods.